# DiViD: Disentangled Video Diffusion for Static–Dynamic Factorization

### v1

@@ 3.2 Inductive Biases for Disentanglement
1. Architectural Bias (Static-Dynamic Residual Encoding)
Inspired by DBSE [3], we adopt the concept of first-frame encoding and frame-residual features. However, we significantly extend this concept by employing an advanced sequence encoder combining convolutional encoders, LSTM, and multi-head self-attention layers. Specifically, the first-frame features $f_{1}$ produce the static token $s$ via a two-layer MLP, while subsequent frames encode residual information $\{f_{i}-f_{1}\}_{i>1}$, processed by LSTM and attention mechanisms to yield frame-specific dynamic tokens. This residual-based encoding inherently discourages static information leakage into dynamic representations.
@@ 4.1 Experimental Setup
Dynamic Encoder. The dynamic encoder models the time-varying content within the sequence. After removing static content by subtraction, the residual vectors are processed by a bidirectional LSTM, followed by a multi-head self-attention block with 8 heads and an embedding dimension of 256. The LSTM output passes through the attention module with a residual connection and layer normalization, then a final projection produces the frame-specific dynamic embeddings.
@@ 4.2.1 Qualitative Results
Source Refer to caption
Target Refer to caption
Ours Refer to caption
DBSE Refer to caption
SPYL Refer to caption
Source Refer to caption
Target Refer to caption
Ours Refer to caption
DBSE Refer to caption
SPYL Refer to caption
@@ 4.2.2 Swap‐based disentanglement evaluation
Model Static Only (%) Dynamic Only (%) Joint Acc. (%) Information Leakage (%)
DBSE 99.0 13.7 13.7 83.5
SPYL 46.1 44.1 17.6 98.5
DiViD 94.1 29.4 28.4 70.9
@@ 4.2.3 Cross‐Leakage Classification
This metric is reported in the last column of Table 1. SPYL shows very high leakage, indicating entangled representations. DBSE still leaks substantially. Our approach reducess average leakage by around 13 percentage points, demonstrating the most effective separation of static appearance and dynamic motion.

### v2

@@ 3.2 Inductive Biases for Disentanglement
1. Architectural Bias (Static–Dynamic Residual Encoding)
Inspired by DBSE [3], we adopt first-frame encoding and frame-residual features. We extend this idea with a Transformer-based sequence encoder—multi-head self-attention with feed-forward blocks, residual connections, and layer normalization. Residuals $r_{i}=f_{i}-f_{1}$ are linearly projected to the model width $z_{\mathrm{dim}}$, augmented with positional encodings $\mathrm{PE}(i)$, and processed by a stack of Transformer encoder layers to yield frame-specific dynamic tokens $d_{i}$. This residual formulation discourages leakage of static information into the dynamic representation.
@@ 3.3 End-to-End Training
Source Refer to caption
Target Refer to caption
Ours Refer to caption
DBSE Refer to caption
SPYL Refer to caption
@@ 4 Experiments
Source Refer to caption
Target Refer to caption
Ours Refer to caption
DBSE Refer to caption
SPYL Refer to caption
@@ 4.1 Experimental Setup
Dynamic Encoder. The dynamic encoder models the time-varying content within the sequence. After removing static content by subtraction, the residual vectors are linearly projected to 256 dimensions, augmented with positional encodings (dropout $0.1$), and processed by a Transformer encoder (4 layers; $d_{\mathrm{model}}=256$; $n_{\text{heads}}=32$; feed-forward size $1024$; dropout $0.1$). The encoder uses residual connections and layer normalization. A final linear head returns 256-dimensional frame-specific dynamic embeddings.
@@ 4.2 Results
Model Static Only (%) $\uparrow$ Dynamic Only (% $\uparrow$) Joint Acc. (%) $\uparrow$ Average Leakage (%) $\downarrow$
DBSE 98.51 16.34 16.34 84.54
SPYL 40.59 45.54 16.83 99.47
DiViD 98.51 31.19 30.20 70.07
@@ 4.2.2 Swap‐based disentanglement evaluation
Method Static Only (%) Dynamic Only (%) Joint Acc. (%)
DiViD w/o Orth 100.0 13.4 13.4
DiViD w/ LSTM 52.0 3.0 1.0
DiViD w/ AdaGN 98.0 24.3 23.8
DiViD w/linear 100.0 22.3 22.3
DiViD 98.5 31.2 30.2
Table 2: Ablation study on MHAD. Joint accuracy requires both identity and action to be correct; marginal accuracies measure each factor independently. Leakage is the average of identity‐into‐motion and motion‐into‐identity leakage rates.
@@ 4.2.3 Cross‐Leakage Classification
This metric is reported in the last column of Table 1. SPYL shows very high leakage, indicating entangled representations. DBSE still leaks substantially. Our approach reducess average leakage by around 14 percentage points, demonstrating the most effective separation of static appearance and dynamic motion.
4.3 Ablation Study
We ablate the main design choices of DiViD on MHAD to quantify the contribution of the inductive biases (Sec. 3.2) and the orthogonality regularizer introduced in our training objective (Sec. 3.3). Specifically, we vary (i) the orthogonality regularizer between static and dynamic tokens, (ii) the temporal encoder used for dynamic residuals, (iii) the conditioning pathway into the U-Net (cross-attention vs. AdaGN), and (iv) the variance ($\beta$) scheduler. Residual subtraction and the shared-noise setting are held fixed across all variants to isolate the effect of each change.
Effect of orthogonality. Removing $\mathcal{L}_{\text{orth}}$ (DiViD w/o Orth) sharply reduces Dynamic-only ($13.4\%$ vs. $31.2\%$ for DiViD) and Joint accuracy ($13.4\%$ vs. $30.2\%$). This confirms that $\mathcal{L}_{\text{orth}}$ is critical to prevent static information from leaking into the dynamic code.
Temporal encoder. Replacing the Transformer encoder with a BiLSTM (DiViD w/ LSTM) collapses dynamic performance (Dynamic-only $3.0\%$, Joint $1.0\%$) under the same training setup, indicating that long-range temporal aggregation and rich token–token interactions from the Transformer are essential for our residual sequence.
Conditioning pathway (U-Net). Substituting cross-attention with AdaGN (DiViD w/ AdaGN) degrades Dynamic-only ($24.3\%$) and Joint ($23.8\%$) relative to DiViD. This supports our claim that cross-attention provides a stronger inductive bias for clean routing of global static $s$ vs. frame-specific $d_{i}$ than global feature modulation.
Variance ($\beta$) scheduler. We examine how the noise variance schedule shapes the time-varying bottleneck in our diffusion decoder. Because we inject the same noise realization across frames, the $\beta$ schedule (and thus the SNR curve) controls how much temporal signal survives at each step, modulating the relative influence of $s$ vs. $d_{i}$ (see the discussion on schedule-dependent bottlenecks in diffusion). DiViD with a cosine schedule (default) attains higher Dynamic-only and Joint accuracy than a linear schedule: from $22.3\%\!\to\!31.2\%$ (Dynamic) and $22.3\%\!\to\!30.2\%$ (Joint), while Static-only remains near saturation ($\approx\!100\%$ vs. $98.5\%$). This indicates that a gentler mid-range SNR decay (cosine) gives cross-attention more opportunity to route motion through $d_{i}$, improving disentanglement, consistent with prior analyses that different $\beta$ schedules induce different information bottlenecks [30].
Summary. DiViD attains the best Joint accuracy ($30.2\%$) by combining: (i) residual subtraction to suppress static leakage at the source, (ii) a Transformer encoder for dynamic tokens, (iii) cross-attention conditioning in the U-Net, (iv) orthogonality regularization, and (v) an appropriate $\beta$ schedule that preserves a useful mid-SNR window for motion routing. Removing $\mathcal{L}_{\text{orth}}$, replacing the Transformer with LSTM/linear encoders, or substituting cross-attention with AdaGN consistently degrades Dynamic-only and Joint metrics under our swap-based protocol (Sec. 4.2).